You have 737 coins that are all either dimes or quarters. If these coins are worth $107, how many of each type of coin do you have?
Found 2 solutions by ewatrrr, nerdybill:
Answer by ewatrrr(24785)


Hi
737 coins
Let x and (737-x) represent the number of quarters and dimes respectively
Question states*** CENTS makes sense
25x + 10(737-x) = 10700 CENTS
Solving for x
15x = 10700 - 7370
x = 222, the number of quarters. 515 the number of dimes. (737-222)
CHECKING our Answer***
25*222 + 10*515 = 10700 or $107.00

Answer by nerdybill(7384)

You have 737 coins that are all either dimes or quarters. If these coins are worth $107, how many of each type of coin do you have?
.
Let q = number of quarters
then
737-q = number of dimes
.
.25q + .10(737-q) = 107
.25q + 73.7-.10q = 107
.15q + 73.7 = 107
.15q = 33.30
q = 222 (number of quarters)
.
number of dimes:
737-q = 737-222 = 515